A parallel-plate capacitor is constructed of two horizontal 12.8-cm-diameter circular plates. A 1.9 g plastic bead with a charge of -5.2 nC is suspended between the two plates by the force of the electric field between them. 1.Which plate, the upper or the lower, is positively charged? 2.What is the charge on the positive plate?
Answer #1

given

d = 12.8 cm

= 0.128 m

m = 1.9 g

= 1.9 x 10-3 kg

C = - 5.2 nC

= - 5.2 x 10-9 C

1 )

the upper plate is positively charged

2 )

using F = mg

F = E q

m g = E q

E = m g / q

= 1.9 x 10-3 x 9.81 / 5.2 x 10-9

E = 3584423.077 N/C

E = / o

= Q / A

= Q / A o

Q = A o E

A = d2 / 4

Q = d2 o E / 4

= 3.14 x 0.1282 x 8.85 x 10-12 x 3584423.077 / 4

Q = 4.08 x 10-7 C

= 40.8 x 10-6 C

so the charge on the positive plate is Q = 40.8 C